OpenStudy (anonymous):

3x^8+6x^7-9x^6+6x^5 what is the greatest common factor & what is the answer help me please

OpenStudy (rajee_sam):

Greatest common factor is the factor that is common to all terms.

OpenStudy (anonymous):

?????

OpenStudy (mathstudent55):

First look at the number part (the coefficients). What is the greatest common factor of 3, 6, -9, and 6? 3 = 3 6 = 2 * 3 -9 = -1 * 3 * 3 6 = 2 * 3 What factor do they all have in common?

OpenStudy (mathstudent55):

Then look at the variable parts: x^8, x^7, x^6, x^5 How many x's do they all have in common?